OpenStudy (anonymous):

mmm, repost: help, TY Sequences: a(sub n)=((3^n)/(4^n+1)), divergent/convergent? if convergent, find the limit

OpenStudy (anonymous):

\[\text{Limit}\left[\frac{3^n}{4^n+1},n\to 0\right]=\frac{1}{2}, \text{Limit}\left[\frac{3^n}{4^n+1},n\to \text{Infinity}\right]=0 \]